XYZ Co. has 9,000 bonds with a price of $925, 15,000 shares of preferred stock with a price of $55, and 250,000 share of common stock with a selling price of $32. Calculate the capital structure weights.
Answer #1

Total value of bonds=(9000*925)=$8,325,000

Total value of preferred stock=(15000*55)=$825,000

Total value of common stock=(250,000*32)=$8,000,000

Total value=$(8,325,000+825000+8,000,000)=$17,150,000

Hence weight of:

Bonds=(8,325,000/17,150,000)=0.4854(or 48.54% approx)

Preferred stock=(825000/17,150,000)=0.0481(or 4.81% approx)

Common stock=(8,000,000/17,150,000)=0.4665 (or 46.65% approx).